Some of you may be familiar with the "Rule of 72". You will need to research this if you are unfamilar with it. This rule is a mathematical construct which has applications in financial planning, population growth, etc..

As a practical exercise, consider whether you can accumulate $1 million (be a millionaire) by merely saving $200 a month.

Furthermore, when you consider this rule within the context of forecasting, describe the importance of choosing your growth rates for projecting your business plans. For example, assume a 3% growth rate verses a 7% growth rate, both on a consistent basis with compounding. Where is your business growth in 20 years under the 2 growth rate assumptions assuming your present day revenue is $1 million?
Answer #1

Introduction:

The Rule of 72 is a mental mathematical construct to estimate the effect of any growth rate, from quick financial calculations to population estimates.

Y = 72 / R and R = 72 / Y

Where Y -years

            r –interest rates

By saving $200 a month results in a total of $2400 yearly. To get the interest rates we divide $2400 by $1million then multiply by 100.we get an interest rates of 0.24%.

To calculate the number of yeas used to double the investment using Rule of 72

Where Y =72/r

We get Y =144

The success business plan depends on a set of targets and objectives incorporated. The overall plan may set strategic goals, may use SMART objectives or targets that is: Specific, Measurable, Achievable, Realistic and Timely.

Targets assist everyone within a business comprehend what they need to achieve and when they need to achieve it. Performance indicators can be used to monitor of employees or a new product. These can be:

  1. Sales or profit amounts over a given period.
  2. Milestones in new product development.
  3. Computation of market-share statistics.

.The formula for annual compound interest is A = P (1 + r/n) ^ nt:

Where:

A = future value of the investment/loan, including interest
P = initial deposit or loan amount
r = the annual interest rate
n = number of times that interest is compounded per year
t = number of years the money is invested or borrowed for

Example 1.

Principal amount: $1000, 000

Annual rate (r):0.03

Compounds per year (n):12

Years (t):20

$1000, 000(1 + 0.03/12) 12(20) = $1820755.00.

Example 2.

Principal amount: $ 1000, 000

Annual amount rate: 0.07

Compounds per year: 12

Years: 20

$1000, 000(1 + 0.07/12) 12(20) = $4038738.85.

References for additional information:

Simons, R. (2013). Levers of control: how managers use innovative control systems to drive strategic renewal. Harvard Business Press.

Allen, D. (2013). Compound Interest-Compound Interest with a Graphing Calculator (SCORM).

Shaw, G., Brown, R., & Bromiley, P. (1998). Strategic stories: How 3M is rewriting business planning. Harvard business review, 76, 41-54.